Solve the following ordinary differential equation.

=> d^2y/dx^2 + dy/dx + y = 0 .
1
Expert's answer
2020-02-18T07:37:13-0500

Solution. The characteristic equation for a ordinary differential equation


"k^2+k+1=0"

Solve the quadratic equation


"D=1^2-4\times 1 \times1=-3<0"

Roots of the equation


"k_1=\frac{-1-\sqrt{3}i}{2}"

"k_2=\frac{-1+\sqrt{3}i}{2}"

Therefore the solution for a ordinary differential equation

"y=e^{\frac{-x}{2}}(C_1sin(\frac{\sqrt{3}x}{2})+C_2cos(\frac{\sqrt{3}x}{2}))"

where C1 and C2 are constants.

Answer.

"y=e^{\frac{-x}{2}}(C_1sin(\frac{\sqrt{3}x}{2})+C_2cos(\frac{\sqrt{3}x}{2}))"

where C1 and C2 are constants.